Max wants to buy a new car in 10 years and estimates that he will need $25,000 to purchase the vehicle of his choice. How much will he have to save annually given that he is in the 25% tax bracket and estimates to earn 6.0% in the market?

A) $1,917 B) $2,745 C) $1,864 D) $2,034 E) $2,304
Answer #1

GIVEN IN THE ABOVE PROBLEM;

MAX NEEDS $ 25,000 IN 10 YEARS TO PURCHASE A CAR

N= 10

i = 6%

PRESENT VALUE OF ORDINARY ANNUITY = C* [1-(1+i) -N] / i

= PV ( 6% , 10YEARS ) =7.3601

= 25000/7.3601

= $3396

THEREFORE AFTER TAX SAVINGS ANNUALLY WOULD BE 3396(1-0.25)= $2547

OPTION B IS THE ANSWER